Can you please help me solve this equation step by step please:

.67 = (1-T)(.8)
Please provide each step in the process. I have forgotten how to do this and have hit a wall.
Thank you.
DRS

Answer by AnlytcPhil(1116)

.67 = (1 - T)(.8)

It might be easier for you if you change
the right sides to

.67 = .8(1 - T)

Now to get rid of decimals, multiply both
sides by 100, which moves the decimals
two places to the right:

(100).67 = (100).8(1 - T)

    67 = 80(1 - T)

    67 = 80 - 80T

Add 80T to both sides

       67 = 80 - 80T
+80T         + 80T
-------------------
 80T + 67 = 80 

Add -67 to both sides

 80T + 67 = 80
     - 67  -67
--------------
 80T     =  13

Divide both sides by 80

     80T =  13 
    ---- = ---- 
     80     80

       T = 13/80

Edwin